# Exact Sequence Interpolation with Transformers

### v2

@@ 3.3. The Transformer
4. Useful Technical Results
In this section, we collect a number of auxiliary technical results that will be instrumental in the proof of our main results. We first revisit the analysis in [2] concerning the asymptotic behavior of tokens evolving under repeated application of hardmax self-attention layers described in (3.1) with $\pi=\pi^{0}$. This asymptotic analysis provides a simplified setting in which we can gain insight into key features of self-attention layers. Beyond being of independent interest as an extension of previous work, the results of this analysis will also inform the design of self-attention layers in the proofs of Theorem˜1.4 and Theorem˜1.6. We then move on to structural lemmas that describe how self-attention layers interact with feed-forward layers inside a transformer block, highlighting geometric constructions that will be used repeatedly in the sequel.
4.1. Asymptotics of Hardmax Self-Attention Layers Revisited
@@
(4.1) $x_{i}(k+1)=\rho x_{i}(k)+\frac{1}{|\mathcal{C}_{i}(X(k),A)|}V\sum_{\ell\in\mathcal{C}_{i}(X(k),A)}x_{\ell}(k),$
where $\mathcal{C}_{i}(X(k),A)$ denotes the hardmax-selected cluster for token $i$ at time $k$. In [2], convergence was established in the special case where
$\rho=1-\gamma,\qquad V=\gamma I_{d},\qquad\gamma\in(0,1),$
with $A\in\mathbb{R}^{d\times d}$ symmetric positive definite. However, the set of possible equilibria in the full-rank positive definite case is extremely rich: it ranges from a fully concentrated state to a fully dispersed one [11, 2]. This motivates us to study additional structural assumptions on $A$ and on the initial configuration to enforce a desired concentration.
Figure 3. Schematic of the transformer architecture described in Section˜$3.$
@@
A natural way to restrict the equilibria is to reduce the rank of $A$. We only consider the symmetric rank-one case
$A=vv^{\top},\qquad v\in\mathbb{R}^{d}\setminus\{0\}.$
@@
Given an initial configuration $\{x_{i}(0)\}_{i\in[n]}$, suppose $v\in\mathbb{R}^{d}$ satisfies $\langle v,x_{i}(0)\rangle\neq 0$ for all $i\in[n]$, and assume there exist unique indices $m,M\in[n]$ such that
@@
Then, the dynamics governed by (4.1) with $\rho=1-\gamma$, $V=\gamma I_{d}$, $\gamma\in(0,1]$, and $A=vv^{\top}$ converge to an equilibrium consisting of $n_{1}$ copies of $x_{m}(0)$ and $n_{2}$ copies of $x_{M}(0)$ where $n_{1}+n_{2}=n$.
@@
Since $A=vv^{\top}$, for every $i,j\in[n]$ we have
$\langle Ax_{i}(k),x_{j}(k)\rangle=\langle v,x_{i}(k)\rangle\,\langle v,x_{j}(k)\rangle.$
At $k=0$, the sets $\mathcal{C}_{i}(X(0),A)$ reduce to
  • a)
    $\mathcal{C}_{i}(X(0),A)=\{M\}$ if $\langle v,x_{i}(0)\rangle>0$,
  • b)
    $\mathcal{C}_{i}(X(0),A)=\{m\}$ if $\langle v,x_{i}(0)\rangle<0$.
We treat $\gamma=1$ separately. In this case, for all $i$ such that $\langle v,x_{i}(0)\rangle>0$, the update rule yields $x_{i}(k)=x_{M}(0)$ for all $k\geq 1$. Similarly for all $i$ such that $\langle v,x_{i}(0)\rangle<0$, $x_{i}(k)=x_{m}(0)$ for all $k\geq 1$.
We now turn to the case $\gamma\in(0,1)$, and show that $a)$ holds for all $k\geq 0$ by induction. Suppose it holds for $k$ and $\langle v,x_{i}(k)\rangle>0$. By $a)$, $\max_{j\in[n]}\langle v,x_{j}(k)\rangle>0$ has a unique maximizer, so
$\mathcal{C}_{i}(X(k),A)=\{M\}.$
Thus the update rule yields
$x_{i}(k+1)=(1-\gamma)x_{i}(k)+\gamma x_{M}(0).$
Taking inner products with $v$ gives
(4.2) $\langle v,x_{i}(k+1)\rangle=(1-\gamma)\langle v,x_{i}(k)\rangle+\gamma\langle v,x_{M}(0)\rangle>0,$
since both terms are positive. Hence $\langle v,x_{i}(k)\rangle$ remains strictly positive for all $k$. Moreover, using (4.2) and that $\gamma\in(0,1)$, we obtain
$\langle v,x_{i}(k+1)\rangle<(1-\gamma)\langle v,x_{M}(0)\rangle+\gamma\langle v,x_{M}(0)\rangle=\langle v,x_{M}(0)\rangle,$
which proves $a)$. An analogous argument shows that $b)$ holds. In particular, maximizing over all $i\in[n]$ yields that $\mathcal{C}_{i}(X(k),A)$ remains equal to $\{M\}$ or $\{m\}$ for all times, depending on the initial sign of $\langle v,x_{i}(0)\rangle$.
It follows that each non-extreme token evolves according to a linear recurrence: if $i$ satisfies $\langle v,x_{i}(0)\rangle>0$, then
$x_{i}(k+1)-x_{M}(0)=(1-\gamma)\big(x_{i}(k)-x_{M}(0)\big).$
Since $\gamma\in(0,1)$, this implies
$x_{i}(k)\to x_{M}(0)\qquad\text{as }k\to\infty.$
Proceeding analogously for $i$ such that $\langle v,x_{i}(0)\rangle<0$ proves the result. ∎
@@
Figure 4. Initial configuration of tokens in the top row, and the asymptotic configuration in the bottom row for $(a)$ Lemma˜4.1, $(b)$ Lemma˜4.2 and $(c)$ Lemma˜4.3.
An illustration of the initial configurations and their asymptotic configurations for Lemmas˜4.1, 4.2 and 4.3 is shown in Figure˜4. Finally, we provide a result to obtain partial clustering with a desired degree of concentration.
@@
Let $A=\xi I_{d}$ with $\xi>0$, $\rho=1-\gamma$, $V=\gamma I_{d}$, $\gamma\in(0,1]$. Let the initial configuration $\{x_{i}(0)\}_{i\in[n]}$ contain pairwise distinct tokens and the subset $\mathcal{I}=\{i^{1},\dots,i^{m}\}\subseteq[n]$ with $|\mathcal{I}|=m$ satisfy
@@
$X^{\star}=\big(\underbrace{x_{i^{1}}(0),\dots,x_{i^{1}}(0)}_{n-m+1\text{ times}},\,x_{i^{2}}(0),\dots,x_{i^{m}}(0)\big),$
@@
Arguing as in the proof of Lemma˜4.2, together with the fact that $x_{i^{j}}(0)\in\mathbb{R}^{d}_{<0}$ for all $j>1$, we obtain $\mathcal{C}_{i}(X(0),A)=\{i^{1}\}$ for all $i\notin\mathcal{I}$, while arguing as in the proof of Lemma˜4.3, we have that $\mathcal{C}_{i}(X(0),A)=\{i\}$ for all $i\in\mathcal{I}$. Thus,
$x_{i}(1)$ $=x_{i}(0)+\gamma(x_{i^{1}}(0)-x_{i}(0))\quad\text{for all}\;\;i\notin\mathcal{I},$
@@
This immediately proves the result when $\gamma=1$. For $\gamma\in(0,1)$, we run an induction argument on $k\geq 1$ as in Lemma˜4.2 to obtain the result. ∎
4.2. Structural Lemmas for Transformer Blocks
We now turn to transformer blocks and investigate how self-attention and feed-forward layers interact, focusing on structural properties that will be essential for proving Theorem˜1.4. Our first question is whether a self-attention layer can leave a specific token from sequence $x_{i}\in X$ unchanged. If $V=(1-\rho)I_{d}$, it is not hard to see from (3.1) that $x_{i}$ will not be changed by a self-attention layer if $\mathcal{C}_{i}(X,A)=\{i\}$. In such a case, we say that the token $x_{i}\in X$ is a leader. The following lemma asserts that a token can be a leader if and only if it is extreme for the closed convex hull of the tokens in the sequence $X$. As usual, we denote by $\operatorname{\overline{co}}(X)$ the closed convex hull of a sequence $X$ and by $\operatorname{ext}\operatorname{\overline{co}}(X)$ the set of its extreme points.
@@ 4.2. Structural Lemmas for Transformer Blocks
(4.3) $\langle Ax_{i},x_{\ell}\rangle<\langle Ax_{i},x_{i}\rangle\quad\forall\ell\neq i.$
@@
(4.4) $\langle Ax_{i},x_{i}\rangle$ $=\langle v,x_{i}\rangle\langle v,x_{i}\rangle$
@@
(4.5) $>-\langle v,x_{i}\rangle\langle v,x_{\ell}\rangle$
(4.6) $=\langle Ax_{i},x_{\ell}\rangle\quad\forall\ell\neq i.\qed$
The next two lemmas establish that all tokens in a sequence $X^{j}$ can be attracted by a single token in that sequence. This is a crucial result for the proof of Theorem˜1.4 and is inspired by the asymptotic dynamics for $A=vv^{\top}$ in Lemma˜4.1. Moreover, the lemmas allow prescribing the attracting token $x_{i^{\star}}^{j^{\star}}$ in a sequence $X^{j^{\star}}$, which is also used in the proof of Theorem˜1.4 iteratively to separate all sequences one by one.
@@
We assume $j^{\star}=1$ without loss of generality, and proceed by induction on $N$. For $N=1$, a vector $v_{1}\in\mathbb{R}^{d}$ such that conditions i)–iii) hold can be chosen using the same arguments in Lemma˜4.5.
Next, we assume that $v_{N-1}\in\mathbb{R}^{d}$ is such that i)–iii) hold for the first N1N-1 sequences, and construct $v_{N}\in\mathbb{R}^{d}$ such that they also hold when the last sequence $X^{N}$ is considered. Because $j^{\star}=1$, only condition ii) can fail when $X^{N}$ is added to the problem. If condition ii) holds also for $j=N$, we can simply set $v_{N}=v_{N-1}$.
@@
(4.7) $\operatorname{\mathrm{FF}}(x_{\ell}^{j})\neq 0\quad\text{and}\quad\mathcal{C}_{\ell}(\mathbb{FF}(X^{j}),vv^{\top})=\{i_{j}\}.$
@@
Choose $v\in\mathbb{R}^{d}$ as in Lemma˜4.6. We now choose the parameters of the FF layer such that ˜4.7 holds. For this, we first observe that adding the same constant vector to every token of every sequence does not change properties ii)–iii) of Lemma˜4.6. We then set the parameters of $\mathbb{FF}$ as $\eta=1$, $W=v/\|v\|^{2}$, $U=0$ and
@@
where $\varepsilon>0$. For this choice, $\sigma(Ux+b)=\sigma(b)=b$, so that $\operatorname{\mathrm{FF}}(x_{\ell}^{j})=x_{\ell}^{j}+b/{\|v\|^{2}}v$. It is clear that, for all $\varepsilon$ sufficiently small, $\operatorname{\mathrm{FF}}(x_{\ell}^{j})\neq 0$ for all $\{\ell,j\}\in[\operatorname{len}(X^{j})]\times[N]$, yielding the first condition in ˜4.7. For the second one, note that
(4.8) $\langle v,\operatorname{\mathrm{FF}}({x}_{\ell}^{j})\rangle$ $=\langle v,x_{\ell}^{j}\rangle+\frac{\langle v,v\rangle}{\|v\|^{2}}b\geq\varepsilon>0$
@@
This fact, combined with property ii) of Lemma˜4.6 implies that $\mathcal{C}_{\ell}(\mathbb{FF}(X^{j}),vv^{\top})=\{i_{j}\}$ for all $\ell\in[\operatorname{len}(X^{j})]$ and all $j\in[N]$. Finally, $i_{j^{\star}}=i^{\star}$ by property iii) of Lemma˜4.6. ∎
Finally, the following lemma is related to the behavior of feed-forward layers and their ability to approximate “hat functions”, moving exactly one token to an arbitrary position, while keeping the rest fixed. This is a distinctive trait of ReLU feed-forward layers and key to arranging tokens in the initial configurations of Lemmas˜4.2, 4.4 and 4.3.
@@
5. Proof of Theorem˜1.4
@@
We now prove Theorem˜1.4 by prescribing explicit transformer parameters. Although the construction is technical, it follows an intuitive strategy that also sheds light on how transformers operate. To emphasize this intuition, we first present the proof strategy in the context of a simple dataset, postponing the full construction to Sections˜5.2, 5.3, 5.4, 5.5 and 5.6.
5.1. Proof Strategy

### v3

@@ 3.3 The transformer
In this work, we fix the $d\times d$ matrix subspaces
$\mathcal{V}=\{\alpha I_{d}:\;\alpha\in\mathbb{R}\},\qquad\mathcal{A}=\{\alpha I_{d}+uv^{\top}:\;\alpha\in\mathbb{R},\;u,v\in\mathbb{R}^{d}\}.$
This implies that our transformer has $\mathcal{O}(d\sum_{j}m^{j})$ parameters. Of course, our construction also works if one replaces these matrix subspaces with larger ones.
4 Useful technical results
In this section, we collect a number of auxiliary technical results that will be instrumental in the proof of our main results. We first revisit the analysis in [alcalde2025clustering] concerning the asymptotic behavior of tokens evolving under repeated application of hardmax self-attention layers described in (3.1) with $\pi=\pi^{0}$. This asymptotic analysis provides a simplified setting in which we can gain insight into key features of self-attention layers. Beyond being of independent interest as an extension of previous work, the results of this analysis will also inform the design of self-attention layers in the proofs of Theorem 1.6 and Theorem 1.8. We then move on to structural lemmas that describe how self-attention layers interact with feed-forward layers inside a transformer block, highlighting geometric constructions that will be used repeatedly in the sequel.
4.1 Asymptotics of hardmax self-attention layers revisited
@@
$x_{i}(k+1)=\rho x_{i}(k)+\frac{1}{|\mathcal{C}_{i}(X(k),A)|}V\sum_{\ell\in\mathcal{C}_{i}(X(k),A)}x_{\ell}(k),$ (4.1)
where $\mathcal{C}_{i}(X(k),A)$ denotes the hardmax-selected cluster for token $i$ at time $k$. In [alcalde2025clustering], convergence was established in the special case where $\rho=1-\gamma$, $V=\gamma I_{d}$, $\gamma\in(0,1)$, with $A\in\mathbb{R}^{d\times d}$ symmetric positive definite. However, the set of possible equilibria in the full-rank positive definite case is extremely rich: it ranges from a fully concentrated state to a fully dispersed one [geshkovski2023emergence, alcalde2025clustering]. This motivates us to study additional structural assumptions on $A$ and on the initial configuration to enforce a desired concentration.
Figure 3: Schematic of the transformer architecture described in Section $3.$
@@
A natural way to restrict the equilibria is to reduce the rank of $A$. We only consider the symmetric rank-one case $A=vv^{\top}$, where $v\in\mathbb{R}^{d}\setminus\{0\}$.
@@
Given an initial configuration $\{x_{i}(0)\}_{i\in[n]}$, suppose $v\in\mathbb{R}^{d}$ satisfies $\langle v,x_{i}(0)\rangle\neq 0$ for all $i\in[n]$, and assume there exist unique indices $m,M\in[n]$ such that $\langle v,x_{M}(0)\rangle>0$, $\langle v,x_{m}(0)\rangle<0$ and
@@
Then, the dynamics governed by (4.1) with $\rho=1-\gamma$, $V=\gamma I_{d}$, $\gamma\in(0,1]$, and $A=vv^{\top}$ converge to an equilibrium consisting of $n_{1}$ copies of $x_{m}(0)$ and $n_{2}$ copies of $x_{M}(0)$ where $n_{1}+n_{2}=n$. For $\gamma=1$, convergence happens in one step.
@@
If $\gamma=1$, the result follows from a direct calculation using (4.1) with $\rho=0$ and $V=I_{d}$. If $\gamma<1$, we start by proving that the following statements hold for all $k\geq 0$:
$\operatorname*{arg\,max}_{j\in[n]}\langle v,x_{j}(k)\rangle=\{M\},$ (4.2)
$\operatorname*{arg\,min}_{j\in[n]}\langle v,x_{j}(k)\rangle=\{m\},$ (4.3)
$x_{M}(k)=x_{M}(0)\quad\text{and}\quad x_{m}(k)=x_{m}(0),$ (4.4)
$\operatorname{sign}(\langle v,x_{i}(k)\rangle)=\operatorname{sign}(\langle v,x_{i}(0)\rangle)\quad\text{for all }i\in[n].$ (4.5)
These properties hold at $k=0$ by the assumptions. We now assume they hold at time $k$ and prove them at time $k+1$. First, we observe that $\langle Ax_{i}(k),x_{j}(k)\rangle=\langle v,x_{i}(k)\rangle\langle v,x_{j}(k)\rangle$, so
$\mathcal{C}_{i}(X(k),A)=\begin{cases}\{M\},&\langle v,x_{i}(k)\rangle>0,\\ \{m\},&\langle v,x_{i}(k)\rangle<0.\end{cases}$
This is because maximizing $\langle v,x_{i}(k)\rangle\langle v,x_{j}(k)\rangle$ over $j$ amounts to maximizing $\langle v,x_{j}(k)\rangle$ if $\langle v,x_{i}(k)\rangle>0$ and minimizing it if $\langle v,x_{i}(k)\rangle<0$, and by the induction hypothesis the unique maximizer and minimizer are $M$ and $m$.
Next, we compute $x_{i}(k+1)$ using the update rule (4.1) with $\rho=1-\gamma$ and $V=\gamma I_{d}$, and take the inner product with $v$ to find that
$\langle v,x_{i}(k+1)\rangle=\begin{cases}(1-\gamma)\langle v,x_{i}(k)\rangle+\gamma\langle v,x_{M}(k)\rangle&\text{if }\langle v,x_{i}(k)\rangle>0,\\ (1-\gamma)\langle v,x_{i}(k)\rangle+\gamma\langle v,x_{m}(k)\rangle&\text{if }\langle v,x_{i}(k)\rangle<0.\end{cases}$ (4.6)
Property 4.5 at time $k$ and our assumptions that $\langle v,x_{m}(0)\rangle<0<\langle v,x_{M}(0)\rangle$ imply that $\langle v,x_{m}(k)\rangle<0<\langle v,x_{M}(k)\rangle$, so we also have $\mathcal{C}_{M}(X(k),A)=\{M\}$ and $\mathcal{C}_{m}(X(k),A)=\{m\}$. Then, since $\gamma\in(0,1]$, we conclude from 4.6 that properties 4.4 and 4.5 hold at time $k+1$. We also conclude that
$\langle v,x_{i}(k+1)\rangle-\langle v,x_{M}(k)\rangle$ $=(1-\gamma)\bigl(\langle v,x_{i}(k)\rangle-\langle v,x_{M}(k)\rangle\bigr)<0$ $\text{if }\langle v,x_{i}(0)\rangle>0,$
$\langle v,x_{i}(k+1)\rangle-\langle v,x_{m}(k)\rangle$ $=(1-\gamma)\bigl(\langle v,x_{i}(k)\rangle-\langle v,x_{m}(k)\rangle\bigr)>0$ $\text{if }\langle v,x_{i}(0)\rangle<0.$
These inequalities imply that $\langle v,x_{m}(k+1)\rangle<\langle v,x_{i}(k+1)\rangle<\langle v,x_{M}(k+1)\rangle$, which in turn imply that properties 4.2 and 4.3 hold at time $k+1$. This concludes the induction process.
Finally, let $n_{1}=\bigl|\{i:\langle v,x_{i}(0)\rangle<0\}\bigr|$, $n_{2}=\bigl|\{i:\langle v,x_{i}(0)\rangle>0\}\bigr|$. For every $i$ with $\langle v,x_{i}(0)\rangle>0$ we have $x_{i}(k+1)-x_{M}(0)=(1-\gamma)\bigl(x_{i}(k)-x_{M}(0)\bigr)$, and therefore
$x_{i}(k)=x_{M}(0)+(1-\gamma)^{k}\bigl(x_{i}(0)-x_{M}(0)\bigr)\to x_{M}(0).$
Similar arguments show that $x_{i}(k)\to x_{m}(0)$ for every $i$ with $\langle v,x_{i}(0)\rangle<0$. Consequently the limiting configuration consists of $n_{1}$ copies of $x_{m}(0)$ and $n_{2}$ copies of $x_{M}(0)$ with $n_{1}+n_{2}=n$. This is easily seen to be an equilibrium for the dynamical map 4.1. ∎
@@
Figure 4: Initial configuration of tokens in the top row, and the asymptotic configuration in the bottom row for $(a)$ Lemma 4.1, $(b)$ Lemma 4.2 and $(c)$ Lemma 4.3.
An illustration of the initial configurations and their asymptotic configurations for Lemmas 4.1, 4.2 and 4.3 is shown in Figure 4. Finally, we provide a result to obtain partial clustering with a desired degree of concentration.
@@
Let $A=\xi I_{d}$ with $\xi>0$, $\rho=1-\gamma$, $V=\gamma I_{d}$, $\gamma\in(0,1]$. Let the initial configuration $\{x_{i}(0)\}_{i\in[n]}$ contain pairwise distinct tokens and the subset $\mathcal{I}=\{i_{1},\dots,i_{m}\}\subseteq[n]$ with $|\mathcal{I}|=m$ satisfy
@@
$X^{\star}=\big(\underbrace{x_{i_{1}}(0),\dots,x_{i_{1}}(0)}_{n-m+1\text{ times}},\,x_{i_{2}}(0),\dots,x_{i_{m}}(0)\big),$
@@
Arguing as in the proof of Lemma 4.2, together with the fact that $x_{i_{j}}(0)\in\mathbb{R}^{d}_{<0}$ for all $j>1$, we obtain $\mathcal{C}_{i}(X(0),A)=\{i_{1}\}$ for all $i\notin\mathcal{I}$, while arguing as in the proof of Lemma 4.3, we have that $\mathcal{C}_{i}(X(0),A)=\{i\}$ for all $i\in\mathcal{I}$. Thus,
$x_{i}(1)$ $=x_{i}(0)+\gamma(x_{i_{1}}(0)-x_{i}(0))\quad\text{for all}\;\;i\notin\mathcal{I},$
@@
This immediately proves the result when $\gamma=1$. For $\gamma\in(0,1)$, we run an induction argument on $k\geq 1$ as in Lemma 4.2 to obtain the result. ∎
4.2 Structural lemmas for transformer blocks
We now turn to transformer blocks and investigate how self-attention and feed-forward layers interact, focusing on structural properties that will be essential for proving Theorem 1.6. Our first question is whether a self-attention layer can leave a specific token $x_{i}$ from a sequence $X^{j}$ unchanged. (Throughout this section, superscripts always refer to sequence indices in the dataset, while subscripts refer to tokens indices within the sequence. In particular, we write $x^{j}_{i}$ to indicate the $i$-th token from sequence $X^{j}$.) If $V=(1-\rho)I_{d}$, it is not hard to see from (3.1) that $x_{i}$ will not be changed by a self-attention layer if $\mathcal{C}_{i}(X,A)=\{i\}$. In such a case, we say that the token $x_{i}\in X$ is a leader. The following lemma asserts that a token can be a leader if and only if it is extreme for the closed convex hull of the tokens in the sequence $X$. As usual, we denote by $\operatorname{\overline{co}}(X)$ the closed convex hull of a sequence $X$ and by $\operatorname{ext}\operatorname{\overline{co}}(X)$ the set of its extreme points.
@@ 4.2 Structural lemmas for transformer blocks
$\langle Ax_{i},x_{\ell}\rangle<\langle Ax_{i},x_{i}\rangle\quad\forall\ell\neq i.$ (4.7)
@@
$\langle Ax_{i},x_{i}\rangle$ $=\langle v,x_{i}\rangle\langle v,x_{i}\rangle$ (4.8)
@@
$>-\langle v,x_{i}\rangle\langle v,x_{\ell}\rangle$ (4.9)
$=\langle Ax_{i},x_{\ell}\rangle\quad\forall\ell\neq i,$ (4.10)
which finishes the proof. ∎
The next two lemmas establish that all tokens in a sequence $X^{j}$ can be attracted by a single token in that sequence. This is a crucial result for the proof of Theorem 1.6 and is inspired by the asymptotic dynamics for $A=vv^{\top}$ in Lemma 4.1. Moreover, the lemmas allow prescribing the attracting token $x_{i^{\star}}^{j^{\star}}$ in a sequence $X^{j^{\star}}$, which is also used in the proof of Theorem 1.6 iteratively to separate all sequences one by one.
@@
We assume $j^{\star}=1$ without loss of generality, and proceed by induction on $N$. For $N=1$, a vector $v_{1}\in\mathbb{R}^{d}$ such that conditions i)–iii) hold can be chosen using the same arguments in Lemma 4.5.
Next, we assume that $v_{N-1}\in\mathbb{R}^{d}$ is such that i)–iii) hold for the first N1N-1 sequences, and construct $v_{N}\in\mathbb{R}^{d}$ such that they also hold when the last sequence $X^{N}$ is considered. Because $j^{\star}=1$, only condition ii) can fail for $v=v_{N-1}$ when $X^{N}$ is added to the problem. If condition ii) holds also for $j=N$, we can simply set $v_{N}=v_{N-1}$.
@@
$\operatorname{\mathrm{FF}}(x_{\ell}^{j})\neq 0\quad\text{and}\quad\mathcal{C}_{\ell}(\mathbb{FF}(X^{j}),vv^{\top})=\{i_{j}\}.$ (4.11)
@@
Choose $v\in\mathbb{R}^{d}$ as in Lemma 4.6. We now choose the parameters of the FF layer such that 4.11 holds. For this, we first observe that adding the same constant vector to every token of every sequence does not change properties ii)–iii) of Lemma 4.6. We then set the parameters of $\mathbb{FF}$ as $\eta=1$, $W=v/\|v\|^{2}$, $U=0$ and
@@
where $\varepsilon>0$. For this choice, $\sigma(Ux+b)=\sigma(b)=b$, so that $\operatorname{\mathrm{FF}}(x_{\ell}^{j})=x_{\ell}^{j}+b/{\|v\|^{2}}v$. It is clear that, for all $\varepsilon$ sufficiently small, $\operatorname{\mathrm{FF}}(x_{\ell}^{j})\neq 0$ for all $\{\ell,j\}\in[\operatorname{len}(X^{j})]\times[N]$, yielding the first condition in 4.11. For the second one, note that
$\langle v,\operatorname{\mathrm{FF}}({x}_{\ell}^{j})\rangle$ $=\langle v,x_{\ell}^{j}\rangle+\frac{\langle v,v\rangle}{\|v\|^{2}}b\geq\varepsilon>0$ (4.12)
@@
This fact, combined with property ii) of Lemma 4.6 implies that $\mathcal{C}_{\ell}(\mathbb{FF}(X^{j}),vv^{\top})=\{i_{j}\}$ for all $\ell\in[\operatorname{len}(X^{j})]$ and all $j\in[N]$. Finally, $i_{j^{\star}}=i^{\star}$ by property iii) of Lemma 4.6. ∎
Finally, the following lemma is related to the behavior of feed-forward layers and their ability to approximate “hat functions”, moving exactly one token to an arbitrary position, while keeping the rest fixed. This is a distinctive trait of ReLU feed-forward layers and key to arranging tokens in the initial configurations of Lemmas 4.2, 4.4 and 4.3.
@@
Finally, we concatenate the two FF layers constructed in Lemmas 4.8 and 4.7 into a single FF layer that implements their composition.
Lemma 4.9.
Let $\operatorname{\mathrm{FF}}^{1}$ and $\operatorname{\mathrm{FF}}^{2}$ be the feed-forward layers from Lemmas 4.8 and 4.7, respectively. There exists a feed-forward layer $\operatorname{\mathrm{FF}}$ of width $d^{\prime}=4$ such that $\operatorname{\mathrm{FF}}=\operatorname{\mathrm{FF}}^{1}\circ\operatorname{\mathrm{FF}}^{2}$.
Proof.
Let $b>0$ and $v$ be chosen as in Lemma 4.7, and set $c=b\|v\|^{-2}v$. The FF layer from Lemma 4.7 satisfies $\operatorname{\mathrm{FF}}^{1}(x)=x+c$. Then, the width-44 FF layer
[…]
satisfies $\operatorname{\mathrm{FF}}(x)=\operatorname{\mathrm{FF}}^{2}(x)+c=\operatorname{\mathrm{FF}}^{1}(\operatorname{\mathrm{FF}}^{2}(x))$, as claimed. ∎
5 Proof of Theorem 1.6
@@
We now prove Theorem 1.6 by prescribing explicit transformer parameters. Although the construction is technical, it follows an intuitive strategy that also sheds light on how transformers operate. To emphasize this intuition, we first present the proof strategy in the context of a simple dataset, postponing the full construction to Sections 5.2, 5.3, 5.4, 5.5 and 5.6.
5.1 Proof strategy